what's the answer to 2 over 9 plus 2 over 9 times 1 over 4 in simplest form
Answer by rfer(16322)

2/9+2/9*1/4=
multiply first
2/9+2/36=
reduce
2/9+1/18=
get common denominator
4/18+1/18=5/18